• 3∧2x - 4(3∧x+1) + 27 = 0
1
Expert's answer
2019-08-06T10:33:19-0400

"3^{2x}-4*3^{x+1}+27=0".

"3^{2x}-12*3^x+27=0."

"3^x=\frac{12-\sqrt{144-108}}{2}\;\;or\;\;3^x=\frac{12+\sqrt{144-108}}{2}"

"3^x=3\;\;or\;\;3^x=9."

"x=1\;\;or\;\;x=2."